OpenStudy (anonymous):

what amout of 1% milk and what amount of 4% milk would you mix to make 5 gallons of 2% milk?

OpenStudy (anonymous):

5 gallons of 2% milk = 0.10 gallons of milk 1 gallon of 1% milk = 0.01 gallons of milk 1 gallon of 4% milk = 0.04 gallons x + y = 5 0.01x + 0.04 y = 0.10 so that simplifies to x + 4y = 10 x + y = 5 So 3y = 5 y = 5/3 or 1.6667 gallons x = 10/3 or 3.333 gallons